# SceneSplat++: A Large Dataset and Comprehensive Benchmark for Language Gaussian Splatting

### v1

@@ 4.2 Key Insights
The primary experimental results from our SceneSplat-Bench benchmark across all selected methods are presented in Tabs. 3 and 4 for indoor scenes and in Tab. 6 for outdoor scenes. The corresponding runtime statistics for individual methods can be found in Tab. 1. SceneSplat – representing the generalizable paradigm – stands out as the clear winner in terms of both performance and efficiency (see Fig. 3 for qualitative examples). Interestingly, in 75% of the experiments, it even outperforms its own pseudo labels used for pretraining (SceneSplat (Pseudo Label)). This result highlights its ability to generalize by leveraging large-scale data priors and learning to predict meaningful vision-language features, effectively mitigating the noise inherent in weak supervision. As illustrated in Fig. 4, visualizations of text query results using enhanced NeRFView tools nerfstudio ; yang2024spectrally ; li2025SceneSplat7k further demonstrate that the generalizable method outperforms the others. Moreover, unlike other methods that rely on extensive feature extraction, the generalizable approach requires only 3DGS as input at inference time and predicts per-primitive vision-language features in a single forward pass.
@@
Among the per-scene methods, optimization-free approaches clearly outperform optimization-based ones, both in terms of segmentation accuracy, as shown in Tabs. 3, 4 and 6, and in runtime efficiency, as reported in Tab. 1. This is likely due to the fact that the objective function in optimization-based methods is typically designed to perform as well as possible on the training views, which may not be ideal to novel viewpoints. Taking this into account, along with the strong performance of the generalizable approach, we conclude that per-scene optimization is not necessary for effective vision-language reasoning in 3DGS. However, within both the optimization-based and optimization-free groups, no single method consistently dominates across all datasets, indicating that performance remains sensitive to dataset-specific characteristics.
@@
A key requirement for reliable benchmarking is having a sufficient number of scenes and appropriately challenging evaluation settings. As shown in Tabs. 3, 4 and 6, the performance of various methods varies noticeably when scaling from a small subset of 10 scenes to the full validation set, highlighting the importance of the benchmark scale in assessing performance. Furthermore,  Tab. 3 demonstrates a clear drop in accuracy when the task complexity increases from 20 to 200 semantic classes, emphasizing the need for demanding benchmarks to reveal the limitations of competing approaches.
@@
Tab. 5 presents our scaling study, which offers two key takeaways. First, training-data scaling consistently improves performance on indoor benchmarks: expanding the training set from 280 to 3503 scenes lifts ScanNet++ f-mIoU from 0.168 $\rightarrow$ 0.284 and ScanNet200 from 0.108 $\rightarrow$ 0.165. A similar effect appears when the model jointly trained on three indoor datasets outperforms that trained on Matterport3D alone. Second—and more surprisingly—an indoor-only model transfers to outdoor scenes, but domain-specific data is still required to close the gap. Without ever seeing outdoor data, the largest indoor model reaches 0.263 mIoU on HoliCity, an improvement of 7.4% over the 280-scene baseline. This suggests that large-scale indoor training induces representations that partially generalize across the indoor–outdoor boundary, see Fig. 5 for zero-shot examples. Nevertheless, a model trained only on HoliCity still leads (0.288 mIoU), indicating that cross-domain capability ultimately needs outdoor supervision. These findings underscore the value of our large-scale indoor–outdoor 3DGS dataset as a powerful training resource.

### v3

@@ 4.2 Key Insights
The primary experimental results from our SceneSplat-Bench benchmark across all selected methods are presented in Tabs.˜3, 4 and 5 for indoor scenes and in Tab.˜6 for outdoor scenes. The corresponding runtime statistics for individual methods can be found in Tab.˜1. SceneSplat – representing the generalizable paradigm – stands out as the clear winner in terms of both performance and efficiency (see Fig.˜3 for qualitative examples). Interestingly, in 75% of the experiments, it even outperforms its own pseudo labels used for pretraining (SceneSplat (Pseudo Label)). This result highlights its ability to generalize by leveraging large-scale data priors and learning to predict meaningful vision-language features, effectively mitigating the noise inherent in weak supervision. In addition, Tab.˜5 shows that the performance on the 2D metric is aligned with the zero-shot performance measured with the 3D metric, with a generalizable method leading the performance. As illustrated in Fig.˜4, visualizations of text query results using enhanced NeRFView tools nerfstudio; yang2024spectrally; li2025SceneSplat7k further demonstrate that the generalizable method outperforms the others. Moreover, unlike other methods that rely on extensive feature extraction, the generalizable approach requires only 3DGS as input at inference time and predicts per-primitive vision-language features in a single forward pass.
@@
Among the per-scene methods, optimization-free approaches clearly outperform optimization-based ones, both in terms of segmentation accuracy, as shown in Tabs.˜3, 4, 6 and 5, and in runtime efficiency, as reported in Tab.˜1. This is likely due to the fact that the objective function in optimization-based methods is typically designed to perform as well as possible on the training views, which may not be ideal to novel viewpoints. Taking this into account, along with the strong performance of the generalizable approach, we conclude that per-scene optimization is not necessary for effective vision-language reasoning in 3DGS. However, within both the optimization-based and optimization-free groups, no single method consistently dominates across all datasets, indicating that performance remains sensitive to dataset-specific characteristics.
@@
A key requirement for reliable benchmarking is having a sufficient number of scenes and appropriately challenging evaluation settings. As shown in Tabs.˜3, 4 and 6, the performance of various methods varies noticeably when scaling from a small subset of 10 scenes to the full validation set, highlighting the importance of the benchmark scale in assessing performance. Furthermore,  Tab.˜3 demonstrates a clear drop in accuracy when the task complexity increases from 20 to 200 semantic classes, emphasizing the need for demanding benchmarks to reveal the limitations of competing approaches.
To further diversify the open-vocabulary evaluation, we include a 3D object localization task that assesses how well each method retrieves objects in 3D space using its constructed 3DGS field. Following the protocol of LangSplat Qin2024LangSplat, we directly select the point with the highest relevance score for each query. Our evaluation is conducted on 2,895 objects from ScanNet and ScanNet++ validation scenes, which is over 12× more objects compared to LERF kerr2023lerf, offering a broader and more realistic assessment. This evaluation is presented in Tab.˜7 and shows similar conclusion as previously discussed evaluations. The generalizable paradigm shows dominant performance, followed by the per-scene optimization-free methods, which outperform per-scene optimization method group.
@@
Tab.˜8 presents our scaling study, which offers two key takeaways. First, training-data scaling consistently improves performance on indoor benchmarks: expanding the training set from 280 to 3503 scenes lifts ScanNet++ f-mIoU from 0.168 $\rightarrow$ 0.284 and ScanNet200 from 0.108 $\rightarrow$ 0.165. A similar effect appears when the model jointly trained on three indoor datasets outperforms that trained on Matterport3D alone. Second—and more surprisingly—an indoor-only model transfers to outdoor scenes, but domain-specific data is still required to close the gap. Without ever seeing outdoor data, the largest indoor model reaches 0.263 mIoU on HoliCity, an improvement of 7.4% over the 280-scene baseline. This suggests that large-scale indoor training induces representations that partially generalize across the indoor–outdoor boundary, see Fig.˜5 for zero-shot examples. Nevertheless, a model trained only on HoliCity still leads (0.288 mIoU), indicating that cross-domain capability ultimately needs outdoor supervision. These findings underscore the value of our large-scale indoor–outdoor 3DGS dataset as a powerful training resource.
@@ 5 Conclusion and Limitations
Acknowledgments and Disclosure of Funding
Yue Li is financially supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy. This work used Dutch national e-infrastructure with the support of the SURF Cooperative under grant no. NWO-2024.035. This work was partially supported by INSAIT, Sofia University “St. Kliment Ohridski” (Partially funded by the Ministry of Education and Science of Bulgaria’s support for INSAIT as part of the Bulgarian National Roadmap for Research Infrastructure), EU Horizon projects ELIAS (No.101120237), and ELLIOT (No.101214398). Some computational resources were provided by the Google Cloud Platform (GCP).